OpenStudy (anonymous):

I REEEEAAALLLY NEED HELP WITH THIS!!! :) What is the equation of the line perpendicular to –x + y = 7 and passing through (-1, -1)?

OpenStudy (anonymous):

-x + y = 2 -x – y = 2 x – y = 2 -x – y = 0 x + y = 0

OpenStudy (misty1212):

lol hi!

OpenStudy (misty1212):

the slope of \(-x+y=7\) is \(1\) the slope of the perpendicular line will be \(-1\)

OpenStudy (anonymous):

hey :)

OpenStudy (misty1212):

equation is \[y+1=-(x+1)\] or \[y+1=-x-1\]

OpenStudy (misty1212):

or \[x+y=-2\]

OpenStudy (misty1212):

or \[-x-y=2\]

OpenStudy (anonymous):

Thank youuuuu ! you are a life saver!